OpenStudy (anonymous):

The product of the page number on two facing pages of a book is 306. Find the page numbers

OpenStudy (anonymous):

17*18=?

OpenStudy (anonymous):

take the square root of 306, it = 17.493. The number below and above it multiply to 306

OpenStudy (anonymous):

x = 1 page x + 1 = the other page x(x + 1) = 306 x^2 + x - 306 = 0 (x - 17)(x + 18) = 0 am I doing this right ehuman ??

OpenStudy (anonymous):

yes you are

OpenStudy (anonymous):

but I am getting a negative number

OpenStudy (anonymous):

I'm trying to figure why one comes up negative though

OpenStudy (anonymous):

I know the answer is 17 and 18, but I can't figure out why my numbers are 17 and -18 ??

OpenStudy (anonymous):

It must be late, I'm baffled

OpenStudy (anonymous):

me too.....my brain hurts

OpenStudy (anonymous):

@Hero ....can you help with this one

OpenStudy (anonymous):

@ranga...we need to borrow your brain

OpenStudy (ranga):

x = 17 or x = -18. Discard the solution that does not apply. A negative page number does not apply. So x = 17 and the next page is x+1 = 18

OpenStudy (anonymous):

ahhhhh x = 17, -18 is extraneous. the answer is x = 17, x+1 = 18

OpenStudy (anonymous):

oh.....I am feeling stupid right now...thanks ranga

hero (hero):

Smart move @ranga

OpenStudy (anonymous):

ranga beat me to the punch while i was typing

OpenStudy (anonymous):

it's lare lol.

OpenStudy (ranga):

Thanks @Hero

OpenStudy (anonymous):

late

OpenStudy (anonymous):

quick fingers

OpenStudy (anonymous):

thanks guys...you always know the answers.